(worth two problems!) A massless spring of spring constant 3800N/m is compressed between two blocks (it
Answer #1

(a)The original potential energy of the system is equal to energy stored in the spring.

If x is compression distance of the spring

PE =1/2kx2 =1/2 (3800 N/m) x2=1900x2

(b)The initial momentum of the system is,

= (1.40 kg) (2.90 m/s)+(2.0 Kg)(2.90 m/s)

= 9.86 kgm/s

(c)Using the conservation of linear momentum along the x axis,

9.86 kg . m/s = (1.40 kg)(ツcos θ)-(2.0 kg)(3.5 m/s)cos 34° w cos θ)-2.897 m/s 0+0-m, (Y sin θ)-m, ( ar sin 34°) Solve the ee